OpenStudy (anonymous):

What is the slope of a line perpendicular to x - 3y = 9?

OpenStudy (anonymous):

@dan815 @shrutipande9

OpenStudy (anonymous):

Convert to slope-intercept form first.

OpenStudy (anonymous):

x - 3y = 9 Subtract 'x' to both sides: -3y = -x + 9 Divide -3 to both sides: y = 1/3x - 3

OpenStudy (amorfide):

get your equation into the standard form y=mx+c then your perpendicular gradient to this line would be the negative reciprocal of m

OpenStudy (anonymous):

Now it's in slope intercept form, y = mx + b, where 'm' is the slope, and 'b' is the y-intercept.

OpenStudy (anonymous):

So what's the slope here?

OpenStudy (anonymous):

idk how to do that stuff

OpenStudy (anonymous):

What's the slope of y = 1/3x - 3? y = mx + b, 'm' is the slope

OpenStudy (anonymous):

so the answer is 1?

OpenStudy (anonymous):

Not quite, it's the fraction before 'x'

OpenStudy (anonymous):

so its 3?

OpenStudy (anonymous):

Fraction. \(\sf y = \dfrac{1}{3}x - 3\)

OpenStudy (anonymous):

The slope is \(\sf\dfrac{1}{3}\).

OpenStudy (anonymous):

Perpendicular lines have opposite reciprocal slopes, so switch the two numbers, can you do that?

OpenStudy (anonymous):

this is confusing!

OpenStudy (anonymous):

Not really. \(\sf y = \color{red}mx + b\) \(\sf y = \color{red}{\dfrac{1}{3}}x-3\) The red parts are the slope.

OpenStudy (anonymous):

no the answer is 1/3?

OpenStudy (anonymous):

so*

OpenStudy (anonymous):

No, that's the slope of this line, it's asking for the slope that's perpendicular to the line.

OpenStudy (anonymous):

so -3?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

thats bullcrap lol

OpenStudy (anonymous):

That's our answer. \(\sf\dfrac{1}{3}\) Flip the fraction: \(\sf\dfrac{3}{1} \rightarrow 3\) Multiply by -1: \(\sf 3 \times -1 = -3\)

OpenStudy (anonymous):

Lol, you'll get used to it :P

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i got like 5 more lol